From: Ian Bearden (bearden@nbi.dk)
Date: Mon Jan 13 2003 - 05:11:09 EST
I have started reducing the data from runs 7341-7350. The reduced data will (I hope!) be written to /brahms/data03/reduced03/seq/... Once this is finished, I'll try to get started on reconstruction and calibrations. Further news as events warrant, Ian -------------------------------------------------------------------- | I.G. Bearden | | Niels Bohr Institute Tb 3 email: bearden@nbi.dk | | Blegdamsvej 17 phone: (+45) 35 32 53 23 | | København Ø FAX: (+45) 31 42 10 16 | | Danmark Www: http://www.nbi.dk/~bearden | -------------------------------------------------------------------- -----Original Message----- From: owner-brahms-l@bnl.gov [mailto:owner-brahms-l@bnl.gov]On Behalf Of Apache Sent: 13. januar 2003 06:41 To: brahms-l@bnl.gov; olchansk@bnl.gov; videbaek@bnl.gov Subject: Shift report 20030113 16:00-24:00 ------------------------------------------------------------ Date: Mon Jan 13 00:40:50 2003, Shift: 16:00-24:00 Supervisor: ykl, with: cr Log book#13 pages 54 - 56 Shift Summary: ------------------------------------------------------------ 1. One store in the afternoon with low d intensity and on long store in the evening with a high d intensity. 2.Reached the goal for 250K trigger 6 at FS setting of 1/8A by run7347. Started FS setting of 1/4A by run7349. SJS and JHL worked on trigger timing and eliminated the noise peak in vertexz spectrum. 3.T3 is still turned off. T1, DVM, and RICH are all working now. ------------------------------------------------------------
This archive was generated by hypermail 2.1.5 : Mon Jan 13 2003 - 05:09:43 EST